Calculate the number of mole presents in certain mass of gas occupying 6.5dm3 at atm and 15•C. ( R=0.082 atm dm3 K-1 mole-1).
1
Expert's answer
2020-01-08T02:02:31-0500

Pressure = 1 atm

temperature = 150C = (273+15)K = 288 K

volume = 6.5 litres

applying ideal gas equation PV=nRT

"\implies" 1*6.5=n*0.082*288

"\implies" n=3.63 moles